ROYAL OPERA HOUSE 12th November: Test

For the record: all feeds were stable at 1107GMT

The feed is now up and stable on IS-905 and E5Wa

The teleport has some issues getting the IS-1002 service working. Please bear with us

This entry was posted in ROH and tagged , , . Bookmark the permalink.